# Leveraging Hallucinations to Reduce Manual Prompt Dependency in Promptable Segmentation

### v1

@@ Abstract
Promptable segmentation typically requires instance-specific manual prompts to guide the segmentation of each desired object. To minimize such a need, task-generic promptable segmentation has been introduced, which employs a single task-generic prompt to segment various images of different objects in the same task. Current methods use Multimodal Large Language Models (MLLMs) to reason detailed instance-specific prompts from a task-generic prompt for improving segmentation accuracy. The effectiveness of this segmentation heavily depends on the precision of these derived prompts. However, MLLMs often suffer hallucinations during reasoning, resulting in inaccurate prompting. While existing methods focus on eliminating hallucinations to improve a model, we argue that MLLM hallucinations can reveal valuable contextual insights when leveraged correctly, as they represent pre-trained large-scale knowledge beyond individual images. In this paper, we utilize hallucinations to mine task-related information from images and verify its accuracy for enhancing precision of the generated prompts. Specifically, we introduce an iterative Prompt-Mask Cycle generation framework (ProMaC) with a prompt generator and a mask generator. The prompt generator uses a multi-scale chain of thought prompting, initially exploring hallucinations for extracting extended contextual knowledge on a test image. These hallucinations are then reduced to formulate precise instance-specific prompts, directing the mask generator to produce masks that are consistent with task semantics by mask semantic alignment. The generated masks iteratively induce the prompt generator to focus more on task-relevant image areas and reduce irrelevant hallucinations, resulting jointly in better prompts and masks. Experiments on 5 benchmarks demonstrate the effectiveness of ProMaC. Code given in https://lwpyh.github.io/ProMaC/.
@@ 4 Experimental Setup
Baselines. To evaluate our approach across various scenarios, we first assess the performance of ProMaC on challenging segmentation tasks, including Camouflaged Object Detection (COD), Medical Image Segmentation (MIS), and Transparent Object Detection (TOD). These tasks are areas where SAM struggle [24]. In the COD task, we compare ProMaC with weakly supervised segmentation methods [27, 60, 58, 60, 17, 17, 20, 22]. Two supervision levels are used for comparison: scribble supervision, where main structures for the foreground and background are sketched during training, and point supervision, where separate points are provided for both foreground and background. In the task-generic prompt setting, we introduce a challenging scenario by only providing a task description as a generic prompt for segmentation. ProMaC integrates LLaVA1.5 [35] with SAM [27]. We also experiment on the MIS and PIS tasks to demonstrate the effectiveness of our method using task-generic prompts compared to previous methods. We assess GPT4V+SAM and LLaVA1.5+SAM in this setting to demonstrate that current MLLM models cannot address it well. We also compare ProMaC against current SOTA promptable segmentation methods to showcase its effectiveness. Next, we evaluate ProMaC on Open-Vocabulary Segmentation (OVS), and compare with leading methods [31, 68, 69, 36, 61, 64, 30]. Our Visual Contrastive Reasoning (VCR) strategy is applicable to other tasks, especially those requiring complex spatial understanding. We used the What’s Up spatial reasoning dataset [25] to evaluate how well VCR guides models to focus on task-relevant regions in Spatial Reasoning (SR) task. Our results are the average of three trials.
@@ 5 Results and Analysis
Visualization. Fig. 4 visually compares our ProMaC with other methods across 3 tasks and also shows the contrastive images we generated. GenSAM handles clear objects well but struggles with complex background. Although GenSAM performs well in complex backgrounds, but struggles with challenging tasks. ProMaC delivers solid segmentation results across different tasks, and our contrastive images remove task-related regions while maintaining semantic and pixel consistency.

### v3

@@ Abstract
Promptable segmentation typically requires instance-specific manual prompts to guide the segmentation of each desired object. To minimize such a need, task-generic promptable segmentation has been introduced, which employs a single task-generic prompt to segment various images of different objects in the same task. Current methods use Multimodal Large Language Models (MLLMs) to reason detailed instance-specific prompts from a task-generic prompt for improving segmentation accuracy. The effectiveness of this segmentation heavily depends on the precision of these derived prompts. However, MLLMs often suffer from hallucinations during reasoning, resulting in inaccurate prompting. While existing methods focus on eliminating hallucinations to improve a model, we argue that MLLM hallucinations can reveal valuable contextual insights when leveraged correctly, as they represent pre-trained large-scale knowledge beyond individual images. In this work, we utilize hallucinations to mine task-related information from images and verify its accuracy for enhancing precision of the generated prompts. Specifically, we introduce an iterative Prompt-Mask Cycle generation framework (ProMaC) with a prompt generator and a mask generator. The prompt generator uses a multi-scale chain of thought prompting, initially exploring hallucinations for extracting extended contextual knowledge on a test image. These hallucinations are then reduced to formulate precise instance-specific prompts, directing the mask generator to produce masks that are consistent with task semantics by mask semantic alignment. The generated masks iteratively induce the prompt generator to focus more on task-relevant image areas and reduce irrelevant hallucinations, jointly resulting in better prompts and masks. Experiments on 5 benchmarks demonstrate the effectiveness of ProMaC. Code given in https://lwpyh.github.io/ProMaC/.
@@ 4 Experimental Setup
Baseline. To evaluate our approach across various scenarios, we first assess the performance of ProMaC on challenging segmentation tasks, including Camouflaged Object Detection (COD), Medical Image Segmentation (MIS), and Transparent Object Detection (TOD). These tasks are areas where SAM struggle [26]. In the COD task, we compare ProMaC with weakly supervised segmentation methods [29, 62, 60, 62, 17, 17, 22, 24, 18]. Two supervision levels are used for comparison: scribble supervision, where main structures for the foreground and background are sketched during training, and point supervision, where separate points are provided for both foreground and background. In the task-generic prompt setting, we introduce a challenging scenario by only providing a task description as a generic prompt for segmentation. ProMaC integrates LLaVA1.5 [37] with SAM [29]. We also experiment on the MIS and PIS tasks to demonstrate the effectiveness of our method using task-generic prompts compared to previous methods. We assess GPT4V+SAM and LLaVA1.5+SAM in this setting to demonstrate that current MLLM models cannot address it well. We also compare ProMaC against current SOTA promptable segmentation methods to showcase its effectiveness. Next, we evaluate ProMaC on Open-Vocabulary Segmentation (OVS), and compare with leading methods [33, 70, 71, 38, 63, 66, 32, 19]. Our Visual Contrastive Reasoning (VCR) strategy is applicable to other tasks, especially those requiring complex spatial understanding. We used the What’s Up spatial reasoning dataset [27] to evaluate how well VCR guides models to focus on task-relevant regions in Spatial Reasoning (SR) task. Our results are the average of three trials.
@@
Figure 5: Visualization of the generated masks and contrastive samples over iterations.
@@ 5 Results and Analysis
Visualization. Fig. 4 and Fig.5 visually compares our ProMaC with other methods across 3 tasks and also shows the contrastive images we generated. GenSAM handles clear objects well but struggles with complex background. Although GenSAM performs well in complex backgrounds, but struggles with challenging tasks. ProMaC delivers solid segmentation results across different tasks, and our contrastive images remove task-related regions while maintaining semantic and pixel consistency.